# Perturbed Decision-Focused Learning for Modeling Strategic Energy Storage

### v1

@@ V-B Energy Storage Behavior Prediction
Evaluation Metric. We employ principles of pattern recognition to assess the efficacy of storage behavior prediction models [34]. Our motivation stems from the unique characteristics of storage operations, which are inherently bi-directional—alternating between charging and discharging phases—and marked by the action’s sparsity, where the output of the storage remains at zero for extended durations. Traditional evaluation metrics, including mean-square error and correlation coefficients, fall short of accurately capturing the prediction quality due to these distinct operational dynamics.
@@ V-B1 Storage Behavior Prediction with Synthetic Data
The historical data of RTP, DAP, and demand data are collected from NYISO within the New York City load zone from 2019 to 2022. These historical data are used as the input features for the proposed and benchmark methods. We use two years for training and one year for testing. The data resolution is one sample per hour. We employ an LSTM network as our prediction layer. The energy storage arbitrage model in (1) generates the energy storage behaviors. The governing reward, i.e., the synthetic price prediction, $\hat{\bm{\lambda}}\in\mathbb{R}^{T}$ is generated by:
@@
We evaluate the performance of our proposed method using energy storage models with two cost terms. The settings for these cost terms are consistent with those described in Section V-A. The hyperparameter $\beta$ is set as $\beta=0.001$. Fig. 3 and Fig. 4 show the prediction performance of the proposed method and two benchmark methods. The ground-truth behaviors and the predictions in four consecutive days are plotted. The proposed method accurately predicts storage behaviors regarding time steps and magnitudes. In contrast, the LSTM method can predict the timing of charge/discharge decisions but struggles to learn power rating information from historical data. Additionally, it mispredicts charge/discharge behaviors during standby periods when energy storage remains inactive. The two-stage method can capture the power rating but predicts a shorter duration. Additionally, it fails to predict energy storage activities on the first two days because the pre-built predictor underestimates the price spread, leading the model to stay on standby.
@@
Figure 5: Energy dispatch of ground-truth, the proposed approach and two benchmark methods: (a) energy storage model with a linear cost term, (b) energy storage model with linear and quadratic cost terms, and (c) energy storage model based on real-world data.
Table III: Comparison of the proposed approach and two benchmark methods, storage model with linear cost term.
@@
Table IV: Comparison of the proposed approach and two benchmark methods, storage model with linear and quadratic cost term.
@@ V-B2 Storage Behavior Prediction with Real-World Data
We conduct experiments on public energy storage data from Queensland University. The Tesla Powerpack battery system has a 1.1 MW power rating and a 2.2 MWh capacity. The battery system is controlled by a demand response engine and engages in two major revenue activities: arbitrage and Frequency Control Ancillary Services (FCAS). According to its performance report [36], the battery system only had a 2-hour duration of FCAS in 2020. Therefore, energy storage behaviors are dominated by storage arbitrage. The data resolution is one sample every 30 minutes. Six months of data are used for training, and four months of data are used for testing. Some days did not have charging/discharging activities due to internet connection issues. Historical energy storage behaviors also have missing data issues for several days. These days are removed from the training and testing data. We only include RTP data as input features.
The parameters of the energy storage arbitrage model are set as follows: power rating $P=1.11MW$, storage efficiency $\eta=0.9$, storage capacity $E=2.2MWh$, the prediction horizon $T=96$, we assume the energy storage model has a linear cost term and a quadratic cost term $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}$, where the coefficients of discharge cost is set as $C_{1}=5$, $C_{2}=3$. The hybrid loss parameter $\beta=0.01$. Note that predicting real-world energy storage behaviors is inherently challenging. First, complete knowledge of model parameters, such as prediction horizon and strategy, is unavailable. Second, our algorithm requires the SoC to be consistent across days, but data quality issues lead to some days being excluded, complicating the learning process. The stochastic nature of real-world energy storage adds further difficulty for accurate behavior predictions. Despite all these challenges, our method consistently outperforms the benchmark methods on this real-world dataset.
@@ VI Conclusions
This paper introduces a decision-focused approach that integrates the physical energy storage model into a neural network-based architecture. Our method incorporates prior knowledge from the storage arbitrage model and infers the hidden reward that guides energy storage decisions. We provide theoretical analysis for the perturbed loss function. To facilitate effective training, we propose a novel hybrid loss function. We demonstrate our proposed framework on two energy storage applications: the self-scheduling energy storage arbitrage problem and the energy storage behavior prediction problem. Numerical results on both applications underscore the significant advantages of our proposed method. In future work, we plan to explore various neural network architectures and select the best-performing networks for our framework. Additionally, we aim to investigate integrating our method into Behind-the-Meter energy disaggregation of energy storage.

### v2

@@ V-B Energy Storage Behavior Prediction
The goal of this experiment is to demonstrate that the proposed decision-focused framework improves the accuracy of predicting energy storage behavior. The LSTM method does not incorporate the energy storage model during training. The two-stage method separates prediction and energy storage optimization into separate parts, training the prediction model solely to minimize the price prediction error without considering the observed charge/discharge behaviors. Our proposed method integrates the energy storage model into the machine learning pipeline, minimizing decision loss with respect to observed charge/discharge actions. We evaluate prediction performance over a full year to demonstrate the effectiveness of our approach.
Evaluation Metric. We employ principles of pattern recognition to assess the efficacy of storage behavior prediction models. Our motivation stems from the unique characteristics of storage operations, which are inherently bi-directional—alternating between charging and discharging phases—and marked by the action’s sparsity, where the output of the storage remains at zero for extended durations. Traditional evaluation metrics, including mean-square error and correlation coefficients, fall short of accurately capturing the prediction quality due to these distinct operational dynamics.
@@ V-B1 Storage Behavior Prediction with Synthetic Data
The historical data for RTP, DAP, and load data are collected from NYISO from 2019 to 2022. We use two years for training and one year for testing. The data resolution is one sample per hour. The dataset collected for this study comprises both input features and target decision values. Similar to the application of energy storage arbitrage, we select typical input features for a price forecaster, including three types: RTP, DAP, and load data. The target decisions are the historical energy storage charge/discharge actions. The input features are structured as a $730\times 24\times 3$ tensor, where 730 represents the number of samples, 24 corresponds to 24 hourly intervals, and 3 denotes the feature types. The target decision data is structured as a $730\times 24$ matrix. The testing dataset follows the same format and contains 365 samples. We employ an LSTM network as our prediction layer. The energy storage arbitrage model in (1) generates the energy storage behaviors. The governing reward, i.e., the synthetic price prediction, $\hat{\bm{\lambda}}\in\mathbb{R}^{T}$ is generated by:
@@
We evaluate the performance of our proposed method using energy storage behaviors with two cost terms. The settings for these cost terms are consistent with those described in Section V-A. The hyperparameter $\beta$ is set as $\beta=0.001$. Fig. 5 and Fig. 7 show the prediction performance of the proposed method and two benchmark methods. The ground-truth behaviors and the predictions in four consecutive days are plotted. The proposed method accurately predicts storage behaviors regarding time steps and magnitudes. In contrast, the LSTM method can predict the timing of charge/discharge decisions but struggles to learn power rating information from historical data. Additionally, it mispredicts charge/discharge behaviors during standby periods when energy storage remains inactive. The two-stage method can capture the power rating but predicts a shorter duration. Additionally, it fails to predict energy storage activities on the first two days. We show the price predictions from the two-stage method with the actual predictions in Fig. 6. As shown in Fig. 6, there are deviations between the two-stage predictions and the actual predictions. While the two-stage method’s predictions follow a similar overall trend to the actual predictions, the two-stage model underestimates the price spread during the first two days. Because the predicted price spread from the two-stage method is lower than the degradation cost, the storage system remains idle.
Figure 6: Price prediction of the two-stage method and the actual prediction. (a) Price prediction of two-stage method in Fig. 5; (b) Price prediction of two-stage method in Fig. 7.
@@
Figure 8: Energy dispatch of ground-truth, the proposed approach and two benchmark methods: energy storage behavior with (a) a linear cost term, (b) linear and quadratic cost terms, and (c) real-world energy storage behavior.
Table III: Comparison of the proposed approach and two benchmark methods, storage behavior with linear cost term.
@@
Table IV: Comparison of the proposed approach and two benchmark methods, storage behavior with linear and quadratic cost term.
@@
Table VI: The performance of the proposed approach without assuming knowledge of the degradation cost, storage behavior with linear cost term and linear and quadratic cost term.
Event-based confusion matrix Magnitude-based confusion matrix
Linear Cost 1840 5116 1435 345 56.18% 79.62% 84.21% 67.40% 1608 5137 1414 577 53.21% 77.21% 73.59% 61.76%
Linear and Quadratic 2197 4751 1082 706 67.00% 79.53% 75.68% 71.08% 1623 4639 1194 1280 57.61% 71.68% 55.91% 56.75%
@@
Note that the regulator can obtain degradation cost information through resource integration registration. Our method does not rely on prior knowledge of linear cost coefficients, and the quadratic cost has only a minor influence on prediction accuracy. The algorithm can interpret both linear and quadratic cost coefficients as part of the reward prediction. However, because the reward prediction is linearly multiplied with charge/discharge decisions in the objective function, the quadratic cost cannot be fully captured, resulting in a slight impact on prediction accuracy. In such cases, the algorithm can leverage data-driven storage model identification techniques [24] to estimate the energy storage degradation quadratic cost term based on inferred pricing dynamics from our approach.
In Fig. 5 and Fig. 7, $C_{1}=10$ for linear case and $C_{1}=5$ and $C_{2}=5$ for linear and quadratic case. In Fig. 10, we do not assume that the degradation cost is known in our optimization model. We set $C_{1}=0$ in our model for the linear cost case. For the linear and quadratic cost case, we set $C_{1}=0$ and $C_{2}=0$. The results, shown in the Fig. 10 and Table VI, indicate that the algorithm achieves similar performance to the case where the ground truth cost terms are assumed. The algorithm accurately predicts the energy storage model under a linear cost. For energy storage with linear and quadratic costs, our method correctly predicts the timing of charge and discharge, with only minor errors in discharge predictions. This demonstrates that our algorithm is not sensitive to prior knowledge of the cost function.
Figure 10: Prediction performance of the proposed method without assuming knowledge of the cost term. Storage behavior with (a) linear cost; (b) linear and quadratic cost.
@@ V-B2 Storage Behavior Prediction with Real-World Data
We conduct experiments on public energy storage data in 2020 from Queensland University. The Tesla Powerpack battery system has a 1.1 MW power rating and a 2.2 MWh capacity. The battery system is controlled by a demand response engine and engages in two major revenue activities: arbitrage and Frequency Control Ancillary Services (FCAS). According to its performance report [39], the battery system only had a 2-hour duration of FCAS in 2020. Therefore, energy storage behaviors are dominated by storage arbitrage. Some days did not have charging/discharging activities due to internet connection issues. Historical energy storage behaviors also have missing data issues for several days. These days are removed from the training and testing data. Six months of data are used for training, and three months of data are used for testing. The data resolution is one sample every 30 minutes. We include RTP data as input features. The input features are structured as a $170\times 96\times 1$ tensor, where 170 represents the number of samples, 96 corresponds to half-hourly intervals, and 1 denotes the feature type. The target decision data is structured as a $170\times 24$ matrix. The testing dataset follows the same format and contains 70 samples.
The parameters of the energy storage arbitrage model are set as follows: power rating $P=1.11MW$, storage efficiency $\eta=0.9$, storage capacity $E=2.2MWh$, the prediction horizon $T=96$, we assume the energy storage behavior has a linear cost term and a quadratic cost term $u(p_{t},b_{t})=C_{1}p_{t}+C_{2}p_{t}^{2}$, where the coefficients of discharge cost is set as $C_{1}=5$, $C_{2}=3$. The hybrid loss parameter $\beta=0.01$. Note that predicting real-world energy storage behaviors is inherently challenging. First, complete knowledge of model parameters, such as prediction horizon and strategy, is unavailable. Second, our algorithm requires the SoC to be consistent across days, but data quality issues lead to some days being excluded, complicating the learning process. The stochastic nature of real-world energy storage adds further difficulty for accurate behavior predictions. Despite all these challenges, our method consistently outperforms the benchmark methods on this real-world dataset.
@@ VI Conclusions
This paper introduces a decision-focused approach that integrates the physical energy storage model into a neural network-based architecture. Our method incorporates prior knowledge from the storage arbitrage model and infers the hidden reward that guides energy storage decisions. We provide theoretical analysis for the perturbed loss function. To facilitate effective training, we propose a novel hybrid loss function. We demonstrate our proposed framework on two energy storage applications: the self-scheduling energy storage arbitrage problem and the energy storage behavior prediction problem. Numerical results on both applications underscore the significant advantages of our proposed method. In future work, we plan to extend our decision-focused framework to strategic energy storage operation under a price-maker scenario, incorporating a grid-based model and accounting for the storage system’s influence on market prices. We will also explore incorporating grid losses [30] into our framework. Additionally, we will integrate our approach into behind-the-meter energy disaggregation for energy storage.